The distance between two telephone poles is
49 m. When a 1.35 kg bird lands on the
telephone wire midway between the poles, the
wire sags 0.185 m.
The acceleration of gravity is 9.8 m/s2 .
How much tension in the wire does the bird
produce? Ignore the weight of the wire.
Answer in units of N.
1
Expert's answer
2010-10-05T10:08:29-0400
T1 | T2 | m*g=0 ⇒ 2*T1*sin (C) =mg ⇒
⇒ T2=T1=mg/(2*AB/AC) = mg/(2*0.185/ √(24.5² + 0.185²)) = 876H